# Joint Velocity-Growth Flow Matching for Single-Cell Dynamics Modeling

### v1

@@ 4.1 Results and Analysis
Ability to reconstruct cellular dynamics. As shown in Tab. 1, our model achieves lowest mean $\mathcal{W}_{1}$ and RME on synthetic datasets, outperforming both unbalancedness-aware methods [25, 22, 49] and unbalancedness-ignoring methods [36, 21], which demonstrate the ability of VGFM to model cellular dynamics for data and mass fitting. We also visualize the learned trajectory and growth function on the Simulation Gene dataset in Figs. LABEL:fig:6pic (a-c). Figures LABEL:fig:6pic (a-c) show that VGFM not only reconstructs accurate trajectories but also produces faithful predictions of spatially varying growth rates.
Scalability and mass-matching performance compared with existing unbalanced models. While existing SOTA approaches of UDSB [49], TIGON [22] and DeepRUOT [25] incorporate unbalanced modeling, these models either introduce biologically implausible transitions [49] (as discussed in [25]) or face difficulties in scaling to high-dimensions [49, 25, 22]. As a result, these methods do not perform well or converge on the 1000-dimensional Gaussian dataset, as in Tab. 1. In contrast, VGFM integrates flow matching based on unbalanced transport, allowing it to scale more effectively to high-dimensional datasets. Our method more accurately recovers the ground-truth dynamics defined in [25] and achieves superior mass-matching accuracy, as shown in Tab. 1 and Figs. LABEL:fig:6pic (d-f).
@@
OT-CFM* [21] 0.302 3.926 10.126
OT-MFM* [36] 0.311 3.976 11.008
UDSB [49] 0.665 0.192 1.914 0.658 25.304 0.965
TIGON [22] 0.099 0.065 1.029 0.542 N/C N/C
DeepRUOT [25] 0.068 0.016 0.474 0.199 N/C N/C
@@
Hold-one-out results on real-world datasets. Table 2 compares hold-one-out results of different methods on EB (5D) and CITE (5D and 50D) datasets. It can be observed that VGFM outperforms most existing approaches in terms of $\mathcal{W}_{1}$ distance. We attribute this to the integration of complementary strengths from both flow matching and simulation-based (Neural ODE) frameworks. On the one hand, flow matching provides a robust initialization for learning the velocity field and growth rate, which helps maintain training stability after incorporating the distribution fitting loss $\mathcal{L}_{\mathrm{OT}}$, leading to better performance than purely simulation-based models, as also shown in Fig. LABEL:fig:eb_vism. On the other hand, since the flow matching objective serves only as an upper bound on the true Wasserstein distance [57], the introduction of $\mathcal{L}_{\mathrm{OT}}$ allows direct optimization of the distribution fitting loss, further enhancing model performance. Moreover, by explicitly modeling cell growth through a time-varying weight function, our framework generalizes beyond the conventional setting where Wasserstein distance is computed between unweighted discrete measures, thereby offering greater flexibility in evaluation and enabling more possible transport solutions.
@@
OT-CFM [21] 0.790 0.882 38.756
SF2M [39] 0.793 0.920 38.524
WLF-UOT [43] 0.738 0.733 37.007
UDSB [49] 1.206 2.023 44.168
OT-MFM [36] 0.713 0.724 36.394
DeepRUOT [25] 0.774 0.648 35.539
VGFM 0.676 0.650 34.182
@@
Comparison on computational efficiency. We compare the computational cost of our method and the SOTA method DeepRUOT [25] in Tab. 3. For fair comparison, all methods in Tab. 3 adopt the same networks 5-layer MLP with 256 hidden units, and are trained on the same GPU until convergence. As shown in Table 3, our full model requires slightly more training time than the variant using only $\mathcal{L}_{\rm VGFM}$, due to the training with additional distribution fitting loss. However, it still achieves faster convergence compared with simulation-based methods of DeepRUOT [25] and the ablation variant without $\mathcal{L}_{\rm VGFM}$. This demonstrates the critical role of the loss $\mathcal{L}_{\rm VGFM}$ for facilitating training convergence.
Table 3: Ablation study on EB (50D) dataset. For comparison, we also report the results of the SOTA approach DeepRUOT [25] in this table.
@@
DeepRUOT [25] 8.169 0.416 9.041 0.415 9.348 0.119 9.808 0.296 90 (mins)
@@ 5 Conclusion, Limitation, and Future Work
The paper proposes the joint Velocity-Growth Flow Matching (VGFM) method that jointly learns state transition and mass growth of single-cell populations via flow matching. VGFM designs an ideal dynamics containing a velocity field and a growth function, driven by our presented two-period dynamic understanding of the static semi-relaxed optimal transport model. Approximating the ideal dynamics using neural networks yields the velocity-growth joint flow matching framework. Extensive experiments on both synthetic and real datasets demonstrate the effectiveness of our method.
@@ A.1 Background: Brenier-Benamou Formulation
We first recall the dynamic formulation of balanced optimal transport, which is known as Brenier-Benamou formula [58]. In this paper, we assume $p_{0}$ and $p_{1}$ are absolutely continuous w.r.t. Lebesgue measure, which is omitted for convenience of description.
@@
$\mathcal{W}_{2}^{2}(p_{0},p_{1})=\inf_{\rho_{t},v_{t}}\left\{\int_{0}^{1}\int_{\Omega}\|v_{t}(\mathbf{x})\|^{2}\mathrm{d}\mathbf{x}\mathrm{d}t\Big{|}\partial_{t}p_{t}=-\nabla\cdot(p_{t}v_{t}),\ p_{t=0}=p_{0},p_{t=1}=p_{1}\right\}$ (A-15)
@@ A.2 Proof of Proposition 1
[…]
By Brenier’s theorem [51], we have $\pi^{*}=(\mathrm{Id},T^{*})_{\#}p_{\lambda}$, where $T^{*}(\mathbf{x})$ is the Monge map from $p_{\lambda}$ to $p_{1}$. Define
@@
Since the second-stage evolution $p_{t}$, for $t\in(\lambda,1]$, is governed solely by the velocity field $v_{t}$ and does not involve mass creation or destruction, both $\tilde{p}_{\lambda}$ and $p_{1}$ have the same total mass. Thus, the Monge map $\tilde{T}^{*}$ from $\tilde{p}_{\lambda}$ to $p_{1}$ under quadratic cost is well-defined. Then, by the Benamou–Brenier formulation [58, 59], we have
@@
$\mathcal{J}_{\rm sot}(\tilde{\pi}^{*})=$ $\mathrm{KL}(\tilde{p}_{\lambda}\|p_{0})+\int_{\Omega^{2}}\|\mathbf{x}_{0}-\mathbf{x}_{1}\|^{2}\mathrm{d}\tilde{\pi}^{*}(\mathbf{x}_{0},\mathbf{x}_{1})$
@@ A.3 Proof and Empirical Evidence of Theorem 1
where $w_{t}$ is the time-dependent weight function.
@@
A.4 Choosing $\tilde{g}_{t}$ According to Eq. (9) Minimizes the $L^{2}$-norm Energy Functional
We aim to prove the following statement: Given a function $w_{t}(\mathbf{x})$ whose logarithmic derivative satisfies $\frac{\mathrm{d}}{\mathrm{d}t}\log w_{t}(\mathbf{x})=g_{t}(\mathbf{x})$ and the boundary condition
@@ B.1 Determine the Entropy Regularization Parameter $\epsilon$ and Relaxation Coefficient $\tau$
Our algorithm involves solving a static semi-relaxed optimal transport problem (11), which includes two critical hyperparameters $\epsilon$ and $\tau$. These parameters play a pivotal role in the behavior and stability of the model. On one hand, if $\epsilon$ is too small, it may lead to numerical instability; if it is too large, it can cause incorrect cross-branch matching in unbalanced data scenarios. To select $\epsilon$, we first set $\tau$ to a moderately large value, such as $\tau=50$, and then perform a grid search over increasing values of $\epsilon$ starting from 0 until numerical stability is achieved. The choice of $\tau$ is even more crucial. When $\tau$ is too small, the KL divergence term between $P_{\#}^{0}\pi$ and $p_{0}$ receives a weak penalty. As a result, the transport mass becomes overly concentrated around points close to $p_{1}$, leading to highly uneven marginals $\pi\mathbf{1}_{m}$ in the discrete setting. This causes instability in the learning of the growth function $g$ and may result in erroneous modeling. Conversely, as $\tau\to+\infty$, Eq. (11) reduces to a balanced OT problem, which contradicts the unbalanced nature of our formulation. To determine a suitable value for $\tau$, we fix the previously selected small $\epsilon$ and gradually increase $\tau$ while observing the variation in $\sum\pi_{ij}c_{ij}$. This curve typically exhibits an increasing-then-flattening trend. Similar to the “elbow metho” used in clustering to select the optimal number of clusters, we identify the region where the curve becomes stable and choose $\tau$ accordingly, as illustrated in Fig. LABEL:fig:tau.
@@ B.2 Simulation of Gene Data
Following the setup of [25], the dynamics of the simulated gene regulatory network are governed by the following system of differential equations:
@@
By setting $\tau=10$ and $\epsilon=0.001$ according to our selection scheme described in Appendix B.1 and illustrated in Fig. LABEL:fig:tau (a), we train the model with a warm-up stage of 200 iterations. After that, distribution fitting loss $\mathcal{L}_{\rm OT}$ is introduced for an additional 30 training epochs. Our model achieve the best performance in terms of (weighted) $\mathcal{W}_{1}$ and relative mass error as shown in Tab. 1.
Table A-5: Simulation parameters on gene regulatory network [25].

### v2

@@ 4.1 Results and Analysis
Ability to reconstruct cellular dynamics. As shown in Tab. 1, our model achieves lowest mean $\mathcal{W}_{1}$ and RME on synthetic datasets, outperforming both unbalancedness-aware methods [zhang2024learning, sha2024reconstructing, pariset2023unbalanced] and unbalancedness-ignoring methods [kapusniak2024metric, tong2024improving], which demonstrate the ability of VGFM to model cellular dynamics for data and mass fitting. We also visualize the learned trajectory and growth function on the Simulation Gene dataset in Figs. 8(a) (a-c). Figs. 8(a) (a-c) show that VGFM not only reconstructs accurate trajectories but also produces faithful predictions of spatially varying growth rates.
Scalability and mass-matching performance compared with existing unbalanced models. While existing SOTA approaches of UDSB [pariset2023unbalanced], TIGON [sha2024reconstructing] and DeepRUOT [zhang2024learning] incorporate unbalanced modeling, these models either introduce biologically implausible transitions [pariset2023unbalanced] (as discussed in [zhang2024learning]) or face difficulties in scaling to high-dimensions [pariset2023unbalanced, zhang2024learning, sha2024reconstructing]. As a result, these methods do not perform well or converge on the 1000-dimensional Gaussian dataset, as in Tab. 1. In contrast, VGFM integrates flow matching based on unbalanced transport, allowing it to scale more effectively to high-dimensional datasets. Our method more accurately recovers the ground-truth dynamics defined in [zhang2024learning] and achieves superior mass-matching accuracy, as shown in Tab. 1 and Figs. 8(a) (d-f).
(a) (a) Predicted dynamics with trajectories of VGFM on Simulation Gene. (b) Predicted and (c) true growth rates on Simulation Gene. Note that for Simulation Gene, the true growth rates could be accessed. (d), (e) and (f) respectively compare predicted relative mass by UDSB, TIGON, DeepRUOT, and VFGM, and observed relative mass from data, on three datasets. Note that in (f), on 50D dataset, the mass predicted by TIGON deviates significantly from the observed trend, because of its difficulty in handling high-dimensional data as discussed in their paper [sha2024reconstructing].
@@
OT-CFM* [tong2024improving] 0.302 3.926 10.126
OT-MFM* [kapusniak2024metric] 0.311 3.976 11.008
UDSB [pariset2023unbalanced] 0.665 0.192 1.914 0.658 N/C N/C
TIGON [sha2024reconstructing] 0.099 0.065 1.029 0.542 N/C N/C
DeepRUOT [zhang2024learning] 0.068 0.016 0.474 0.199 N/C N/C
@@
Hold-one-out results on real-world datasets. Table 2 compares hold-one-out results of different methods on EB (5D) and CITE (5D and 50D) datasets. It can be observed that VGFM outperforms most existing approaches in terms of $\mathcal{W}_{1}$ distance. We attribute this to the integration of complementary strengths from both flow matching and simulation-based (Neural ODE) frameworks. On the one hand, flow matching provides a robust initialization for learning the velocity field and growth rate, which helps maintain training stability after incorporating the distribution fitting loss $\mathcal{L}_{\mathrm{OT}}$, leading to better performance than purely simulation-based models, as also shown in Fig. 11(a). On the other hand, since the flow matching objective serves only as an upper bound on the true Wasserstein distance [benton2024error], the introduction of $\mathcal{L}_{\mathrm{OT}}$ allows direct optimization of the distribution fitting loss, further enhancing model performance. Moreover, by explicitly modeling cell growth through a time-varying weight function, our framework generalizes beyond the conventional setting where Wasserstein distance is computed between unweighted discrete measures, thereby offering greater flexibility in evaluation and enabling more possible transport solutions.
Pancreas dataset under 2000-dimensional gene space. To further explore the scalability of VGFM, we applied our model and compared methods explicitly modeling $g_{t}(x)$ [sha2024reconstructing, zhang2024learning] to the Pancreas dataset with 2000 highly variable genes. We first observed that VGFM is the only method showing a steadily decreasing training loss, both for $\mathcal{L}_{\mathrm{VGFM}}$ and $\mathcal{L}_{\mathrm{OT}}$, as shown in Fig.A-11. Next, following [palma2025multi], we calculated the means and variances of the real and generated gene at day 15.5 and plotted the corresponding mean-variance trend and histograms. The results show that the generated samples closely follow the real data, as shown in Fig. 24(a). Finally, we analyzed the learned $g_{w}(x,t)$ and found that our model successfully reconstructs the unbalanced pattern and identifies key genes without being given any prior knowledge of the cell types. The results are reported in Fig. A-24, A-25 and 27(a). For more details, please refer to Appendix B.8.
@@
OT-CFM [tong2024improving] 0.790 0.882 38.756
SF2M [tong2024simulation] 0.793 0.920 38.524
UDSB [pariset2023unbalanced] 1.206 2.023 44.168
OT-MFM [kapusniak2024metric] 0.713 0.724 36.394
DeepRUOT [zhang2024learning] 0.774 0.845 38.681
VGFM 0.676 0.745 37.386
@@
(a) Visualization of predicted dynamics by (a) VFGM (w/o $\mathcal{L}_{\rm OT}$) and (b) VGFM, and growth rate by (c) VFGM, on EB (5D) dataset, where the hold-out time is the first intermediate timepoint.
Comparison on computational efficiency. We compare the computational cost of our method and the SOTA method DeepRUOT [zhang2024learning] in Tab. 3. For fair comparison, all methods in Tab. 3 adopt the same networks 5-layer MLP with 256 hidden units, and are trained on the same GPU until convergence. As shown in Table 3, our full model requires slightly more training time than the variant using only $\mathcal{L}_{\rm VGFM}$, due to the training with additional distribution fitting loss. However, it still achieves faster convergence compared with simulation-based methods of DeepRUOT [zhang2024learning] and the ablation variant without $\mathcal{L}_{\rm VGFM}$. This demonstrates the critical role of the loss $\mathcal{L}_{\rm VGFM}$ for facilitating training convergence.
To make $\mathcal{L}_{\mathrm{OT}}$ a fully differentiable component in our training pipeline, we also employed the Sinkhorn divergence [feydy19a] to replace $\mathcal{W}_{1}$ as the distribution fitting loss $\mathcal{L}_{\mathrm{OT}}$, forming a variant of VGFM denoted as VGFM(*) in Tab. 3. It can be observed that using the Sinkhorn divergence leads to a further improvement in the accuracy of dynamics reconstruction, while reducing the computational cost for distribution fitting. For more details, please refer to Appendix B.6.
Table 3: Ablation study on EB (50D) dataset. For comparison, we also report the results of the SOTA approach DeepRUOT [zhang2024learning] in this table.
@@
DeepRUOT [zhang2024learning] 8.169 0.416 9.041 0.415 9.348 0.119 9.808 0.296 90 (mins)
@@ 5 Conclusion, Limitation, and Future Work
The paper proposes the joint Velocity-Growth Flow Matching (VGFM) method that jointly learns state transition and mass growth of single-cell populations via flow matching. VGFM designs an ideal dynamics containing a velocity field and a growth function, driven by our presented two-period dynamic understanding of the static semi-relaxed optimal transport model. Approximating the ideal dynamics using neural networks yields the velocity-growth joint flow matching framework.
@@
Acknowledgment
This work was supported by the National Key R&D Program 2021YFA1003002, NSFC (12125104, 623B2084, 12501709, 12426313), China National Postdoctoral Program for Innovative Talents (BX20240276), China Fundamental Research Funds for the Central Universities (xzy022025047) and China Postdoctoral Science Foundation (2025M773058).
Outline of Appendix
The appendix is organized into three main parts.
Appendix A presents the theoretical foundations. We begin by reviewing the Brenier–Benamou theorem [benamou2000computational], followed by detailed proofs of Proposition 1 and Theorem 1 that are stated in the main text, as well as the justification for our choice of the reparameterized growth function $\tilde{g}$ in Eq. (9).
Appendix B focuses on experiments. We first describe our general implementation setup, including the strategy for selecting the hyperparameters in Eq. (11). Next, we provide detailed settings, visualizations and analysis for each experiment.
@@ A.1 Background: Brenier-Benamou Formulation
We first recall the dynamic formulation of balanced optimal transport, which is known as Brenier-Benamou formula [benamou2000computational]. In this paper, we assume $p_{0}$ and $p_{1}$ are absolutely continuous w.r.t. Lebesgue measure, which is omitted for convenience of description.
@@
$\mathcal{W}_{2}^{2}(p_{0},p_{1})=\inf_{p_{t},v_{t}}\left\{\int_{0}^{1}\int_{\Omega}\|v_{t}(\mathbf{x})\|^{2}p_{t}(\mathbf{x})\ \mathrm{d}\mathbf{x}\mathrm{d}t\Big|\partial_{t}p_{t}=-\nabla\cdot(p_{t}v_{t}),\ p_{t=0}=p_{0},p_{t=1}=p_{1}\right\}$ (A-15)
@@ A.2 Proof of Proposition 1
This implies the continuity equation $\partial_{t}p_{t}=g_{t}p_{t}$ for $t\in[0,\lambda]$, and hence we have $p_{0}(\mathbf{x})\exp\left(\int_{0}^{\lambda}g_{t}(\mathbf{x})dt\right)={\rm P}^{0}_{\#}\pi^{*}$.
By Brenier’s theorem [peyre_computational_2020], we have $\pi^{*}=(\mathrm{Id},T^{*})_{\#}p_{\lambda}$, where $T^{*}(\mathbf{x})$ is the Monge map from $p_{\lambda}$ to $p_{1}$. Define
@@
Since the second-stage evolution $p_{t}$, for $t\in(\lambda,1]$, is governed solely by the velocity field $v_{t}$ and does not involve mass creation or destruction, both $\tilde{p}_{\lambda}$ and $p_{1}$ have the same total mass. Thus, the Monge map $\tilde{T}^{*}$ from $\tilde{p}_{\lambda}$ to $p_{1}$ under quadratic cost is well-defined. Then, by the Benamou–Brenier formulation [benamou2000computational, figalli2021invitation], we have
@@
[…]
@@ A.3 Proof and Empirical Evidence of Theorem 1
where $w_{t}$ is the time-dependent weight function and $\mathbb{I}_{\Omega}$ is the indicator function, for any function $f,f\cdot\mathbb{I}_{\Omega}(t)=\begin{cases}f,\quad\text{If $t\in\Omega$,}\\ 0,\quad\text{Otherwise.}\end{cases}$.
@@
A.4 Motivation for time-independent growth function $\tilde{g}_{t}$ according to Eq. (9)
Minimization of $L^{2}$ energy potential. We aim to prove the following statement: Given a function $w_{t}(\mathbf{x})$ whose logarithmic derivative satisfies $\frac{\mathrm{d}}{\mathrm{d}t}\log w_{t}(\mathbf{x})=g_{t}(\mathbf{x})$ and the boundary condition
@@
Explanation from the Malthusian growth model [Malthus1798], i.e., the exponential growth model. The Malthusian growth model is $\frac{\mathrm{d}p_{t}}{\mathrm{d}t}=gp_{t}$, where $p_{t}$ is the population size and $g$ is a constant growth rate. Our model on the growth $\frac{\partial p_{t}}{\partial t}=g_{t}p_{t}$ is consistent with the above Malthusian growth model. In these models, $g$ is treated as a constant based on the assumption that the resources are abundant and the environment is stable, causing growth rates to remain relatively stable over time. In the context of scRNA-seq experiments, these conditions are typically satisfied because cells are often cultured or sampled under controlled laboratory conditions, where nutrient supply, temperature, and other environmental factors are maintained at constant levels. Thus, choosing this form is biologically reasonable, especially when there is no prior knowledge about the growth rate.
@@ B.1 Determine the Entropy Regularization Parameter $\epsilon$ and Relaxation Coefficient $\tau$
Our algorithm involves solving a static semi-relaxed optimal transport problem (11), which includes two critical hyperparameters $\epsilon$ and $\tau$. These parameters play a pivotal role in the behavior and stability of the model. On one hand, if $\epsilon$ is too small, it may lead to numerical instability; if it is too large, it can cause incorrect cross-branch matching in unbalanced data scenarios. To select $\epsilon$, we first set $\tau$ to a moderately large value, such as $\tau=50$, and then perform a grid search over increasing values of $\epsilon$ starting from 0 until numerical stability is achieved. The choice of $\tau$ is even more crucial. When $\tau$ is too small, the KL divergence term between $P_{\#}^{0}\pi$ and $p_{0}$ receives a weak penalty. As a result, the transport mass becomes overly concentrated around points close to $p_{1}$, leading to highly uneven marginals $\pi\mathbf{1}_{m}$ in the discrete setting. This causes instability in the learning of the growth function $g$ and may result in erroneous modeling. Conversely, as $\tau\to+\infty$, Eq. (11) reduces to a balanced OT problem, which contradicts the unbalanced nature of our formulation. To determine a suitable value for $\tau$, we fix the previously selected small $\epsilon$ and gradually increase $\tau$ while observing the variation in $\sum\pi_{ij}c_{ij}$. This curve typically exhibits an increasing-then-flattening trend. Similar to the “elbow metho” used in clustering to select the optimal number of clusters, we identify the region where the curve becomes stable and choose $\tau$ accordingly, as illustrated in Fig. 14(a).
@@
(a) Fixing $\epsilon$, we plot the transport cost as a function of $\tau$. (a) In Simulation Gene, the curve flattens around $\tau=5$; (b) In Dyngen, the transport cost also stabilizes near $\tau=5$.
@@ B.2 Simulation of Gene Data
Following the setup of [zhang2024learning], the dynamics of the simulated gene regulatory network are governed by the following system of differential equations:
@@
By setting $\tau=10$ and $\epsilon=0.003$ according to our selection scheme described in Appendix B.1 and illustrated in Fig. 14(a) (a). Our model achieve the best performance in terms of (weighted) $\mathcal{W}_{1}$ and relative mass error as shown in Tab. 1.
Table A-5: Simulation parameters on gene regulatory network [zhang2024learning].
@@
Growth rate correlation analysis.
To validate the accuracy of our growth rate modeling and learning, we perform correlation analysis between the predicted growth rates and the ground truth. We utilize this dataset, which is generated from the prescribed dynamical systems (Eq. (A-23). Since the ground truth growth rates can be computed analytically, we conduct correlation analysis on out-of-distribution (OOD) time points—excluded from the training data—to assess the generalization capability of VGFM. The results are summarized in Table A-6.
Table A-6: Correlation analysis on four OOD time points for Simulation Gene Data
OOD Time Point Pearson Correlation Coefficient
$t_{0.5}$ 0.980
$t_{1.5}$ 0.992
$t_{2.5}$ 0.995
$t_{3.5}$ 0.996
The growth rates predicted by our VGFM exhibit strong correlation with the ground truth values.
Uncertainty quantification. Although our current model is deterministic, we incorporate dropout into the velocity and growth networks to quantify predictive uncertainty. During inference, we compute the average variance for each dimension of velocity and growth across all data points, considering different branches (0 denotes the quiescent region i.e., the lower left corner of Fig. 8(a) while 1 denotes the region where exhibits mass variation and state transition) and time points. We conduct the experiment with a dropout rate of 0.1, performing 5 stochastic forward passes through the velocity and growth networks for all data points. The results are presented in Table A-7.
Table A-7: Uncertainty estimation of predicted velocity (2-dimensional) and growth on Simulation Gene Data. $\text{var}_{x_{1}}$ and $\text{var}_{x_{2}}$ represent the average variance of the first and second velocity dimensions, respectively, while $\text{var}_{g}$ denotes the variance of growth. All variance values are scaled by $10^{-3}$.
Branch Time $\mathbf{\text{var}_{x_{1}}}$ ($\times 10^{-3}$) $\mathbf{\text{var}_{x_{2}}}$ ($\times 10^{-3}$) $\mathbf{\text{var}_{g}}$ ($\times 10^{-3}$)
0 0 0.1080 0.1000 0.0550
0 1 0.0770 0.0600 0.0660
0 2 0.0660 0.0670 0.1100
0 3 0.0850 0.0820 0.1910
0 4 0.1220 0.1370 0.3500
1 0 2.9320 6.5010 0.2510
1 1 3.0230 4.2870 0.3060
1 2 2.0410 1.9780 0.5590
1 3 0.9530 0.9760 0.7660
1 4 0.2800 0.3100 1.1680
We analyze the uncertainty patterns from both spatial and temporal perspectives:
For spatial analysis, cells in branch 0 exhibit minimal movement with little variation in state and counts. Correspondingly, the variances of $v_{1}$, $v_{2}$, and $g$ in this branch are consistently small, indicating low predictive uncertainty. In contrast, branch 1 demonstrates substantially larger variances, reflecting higher uncertainty in predictions.
For temporal analysis, The Simulation Gene Data features significant state transitions and quantity changes during initial phases, with diminishing magnitudes over time. The variance patterns of $v_{1}$, $v_{2}$, and $g$ accurately capture this temporal evolution, validating our uncertainty estimation approach. Specifically, for branch 1, the of $v$ decrease monotonically over time, aligning with the reduced dynamics in later stages.
Effect of sampling bias. In cellular dynamics reconstruction, it is conventionally assumed that observed datasets accurately capture the temporal evolution of the true cell distribution. Consequently, the observed cell counts should align with the biological processes of cell proliferation and death. Evaluating the robustness of VGFM when trained on biased observational data holds significant practical relevance.
To generate non-uniform or subsampled data, we resample the original dataset with perturbed branch ratios. Specifically, for each time point $t$, we randomly sample a perturbation factor $a_{t}\sim\mathcal{N}(0,\sigma^{2})$. Given an original branch ratio of $s_{t}:(1-s_{t})$, the resampled ratio becomes $(s_{t}-|a_{t}|):(1-s_{t}+|a_{t}|)$, where $\sigma$ quantifies the bias level. We train our model on the resampled data and evaluate its performance on the original dataset to assess VGFM’s robustness.
Table A-8: Performance evaluation ($\mathcal{W}_{1}$/RME metrics) on Simulation Gene Data with varying sampling bias levels, averaged over three random seeds.
Bias Factor $\sigma$ $t_{1}$ $t_{2}$ $t_{3}$ $t_{4}$
0 (unbiased) 0.041/0.007 0.053/0.005 0.038/0.007 0.040/0.008
0.1 0.074/0.035 0.094/0.051 0.075/0.063 0.077/0.032
0.2 0.085/0.059 0.113/0.073 0.134/0.062 0.123/0.125